Determine the molecular formula of alaxin C5H8nO5 if it's relative molecular Mass is 284

1
Expert's answer
2022-06-07T19:03:04-0400

(Empirical formula) n = Molecular formula

(C5H8)nO5 = 284 ......... (I)

[(12*5) + (1*8)]n + (16*5) = 284

(60 + 8)n + 80 = 284

68n + 80 = 284

68n = 284 - 80 = 204

n = 204/68 = 3


Substitute for n = 3 in equation (I)

(C5H8)3O5

C15H24O5